OpenStudy (anonymous):

1) Write an equation in point-slope form for the line through the given point with the given slope. (5, 2); m = 3 2) Write an equation in point-slope form for the line through the given point with the given slope. (-10, -6); m = -5/8 3) Write y = (-3/4)x + 3 in standard form using integers.

pooja195 (pooja195):

For 1 and 2 use this \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

2 - 2 = 3( 5 - 5)?

pooja195 (pooja195):

O_o 1) x1=5 y1=2 m=3 plug them into the formula y-y1=m(x-x1)

OpenStudy (anonymous):

y - 2 = 3( x - 5)

pooja195 (pooja195):

perfect! Do the same for #2

OpenStudy (anonymous):

2) x1 = -10 y1 = -6 m = -5/8 y - y1 = m(x - x1) y - (-6) = -5/8(x - (-10)

pooja195 (pooja195):

perfect :D BUT when we have 2 negative signs next to each other they become postitive :) \[\huge~\rm~y +6 = -5/8(x +10) \]

OpenStudy (anonymous):

Thanks!

pooja195 (pooja195):

=) Ok i am not sure about #3 i havent done that in a while perhaps @Photon336 can explain that part?

OpenStudy (anonymous):

okay

pooja195 (pooja195):

@Luigi0210

OpenStudy (luigi0210):

Do you know what standard form is?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

Basically, it wants you to put it in standard form without any fractions/decimals. I'm assuming this is "Standard Form of A Linear Equation" so we use \(\large Ax+By=C \) Where A cannot be negative and A&B shouldn't be 0, and A, B, and C are integers (:

OpenStudy (anonymous):

So like 3x - 4y = 12?

OpenStudy (luigi0210):

Yup! Except that -4y should be a +4y (: \(\large y=-\frac{3}{4}x+3 \) \(\large \frac{3}{4} x +y=3 \) \(\large 3x+4y=12 \)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (luigi0210):

No problem!

pooja195 (pooja195):

Good job bui!! :D